# MEGA: Memory-Efficient 4D Gaussian Splatting for Dynamic Scenes

### v1

@@ Abstract
4D Gaussian Splatting (4DGS) has recently emerged as a promising technique for capturing complex dynamic 3D scenes with high fidelity. It utilizes a 4D Gaussian representation and a GPU-friendly rasterizer, enabling rapid rendering speeds. Despite its advantages, 4DGS faces significant challenges, notably the requirement of millions of 4D Gaussians, each with extensive associated attributes, leading to substantial memory and storage cost. This paper introduces a memory-efficient framework for 4DGS. We streamline the color attribute by decomposing it into a per-Gaussian direct color component with only 3 parameters and a shared lightweight alternating current color predictor. This approach eliminates the need for spherical harmonics coefficients, which typically involve up to 144 parameters in classic 4DGS, thereby creating a memory-efficient 4D Gaussian representation. Furthermore, we introduce an entropy-constrained Gaussian deformation technique that uses a deformation field to expand the action range of each Gaussian and integrates an opacity-based entropy loss to limit the number of Gaussians, thus forcing our model to use as few Gaussians as possible to fit a dynamic scene well. With simple half-precision storage and zip compression, our framework achieves a storage reduction by approximately 190$\times$ and 125$\times$ on the Technicolor and Neural 3D Video datasets, respectively, compared to the original 4DGS. Meanwhile, it maintains comparable rendering speeds and scene representation quality, setting a new standard in the field.
Figure 1: Our approach significantly reduces storage requirements while maintaining comparable photorealistic quality and real-time rendering speed with 4D Gaussian Splatting (4DGS) (Yang et al., 2024a). The core idea is to develop a memory-efficient 4D Gaussian representation and use as few Gaussians as possible to fit dynamic scenes well. (a) 4DGS requires up to 13 million Gaussians to render the Birthday scene, whereas our method only needs 0.91 million Gaussians. (b) Quantitative comparisons of rendering quality, storage size, and speed against various competitive baselines on the Technicolor dataset.
@@ 1 Introduction
Dynamic scene reconstruction from multi-view videos is gaining widespread interest in computer vision and graphics due to its broad applications in virtual reality (VR), augmented reality (AR), and 3D content production. The emergence of neural radiance field (NeRF) (Mildenhall et al., 2021) enables high-quality novel view synthesis from multi-view image inputs. It has been further extended to represent dynamic scenes by modeling a direct mapping from spatio-temporal coordinates to color and density (Pumarola et al., 2021; Li et al., 2022b; Cao & Johnson, 2023). Despite the impressive visual quality of NeRF-based methods, they require dense sampling along rays, leading to slow rendering speeds that hinder practical applications.
The recent introduction of 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) marks a significant shift in the field of novel view synthesis. This approach incorporates the explicit 3D Gaussian representation and differentiable tile-based rasterization to enable real-time rendering speeds that significantly outperform NeRF-based methods. Built on this framework, subsequent studies have developed 4D Gaussian Splatting (4DGS) (Yang et al., 2024a; Duan et al., 2024), which conceptualizes scene variations across different timestamps as 4D spatio-temporal Gaussian hyper-cylinder. As shown in Fig. 2, when depicting a 3D scene at a given timestamp, these 4D Gaussians will first be sliced into 3D Gaussians with time-varying positions and opacity. Then, the 3D Gaussians with the temporal decay opacity below a specific threshold are filtered out. This filtering operation helps 4DGS to describe the transient content such as emerging or vanishing objects. Finally, following 3DGS, the remaining 3D Gaussians are projected onto 2D screens through fast rasterization. By directly optimizing a collection of 4D Gaussians, 4DGS effectively captures both static and dynamic scene elements, thereby achieving photorealistic visual quality.
@@
Nevertheless, compacting the properties of the 4D Gaussian alone cannot effectively alleviate the problem of excessive number of Gaussians required. Existing 4DGS baselines (Yang et al., 2024a; Duan et al., 2024) assume that each sliced 4D Gaussian exhibits only linear movement over time while maintaining constant covariance, which means that the complex motion in the scene has to be modeled by a combination of multiple Gaussians. Moreover, as illustrated in Fig. 4 (a), only about 6% of Gaussians actively participate in rendering at any given time, because the temporal decay opacity forces each Gaussian to be visible only near its mean time center and invisible at other times. These inherent properties significantly limit the effective utilization of each Gaussian, thereby increasing the number of Gaussians needed for adequate scene rendering. To overcome this limitation, we introduce an efficient entropy-constrained Gaussian deformation field designed to expand the operational range of 4D Gaussians. This deformation model leverages both temporal and viewpoint information to accurately represent Gaussian motion, shape, and transience changes. Meanwhile, a spatial opacity-based entropy loss is introduced to push the spatial opacity of each Gaussian towards binary states (either one or zero). This adjustment aids in identifying and eliminating non-essential Gaussians that contribute minimally to the overall performance. In this way, our proposed strategy not only effectively reduces the number of Gaussians, but also improves the utilization rate of the Gaussians involved in rendering given the time and viewing angle. Finally, to store the parameters of our streamlined 4DGS, we employ 16-bit floating-point (FP16) precision with zip delta compression algorithm to achieve further reductions in memory footprint. In summary, our main contributions are three-fold:
@@ 2 Related Works
Neural Rendering for Static Scenes. Recently, the advent of neural rendering has attracted increasing interest in 3D scene representation and reconstruction. NeRF, pioneered by Mildenhall et al. (2021), represents the volume density and view-dependent emitted radiance of a 3D scene as a function of 5D coordinates (3D position and 2D viewing direction) using an MLP. However, the vanilla NeRF relies solely on a large MLP to store scene information, significantly limiting its training and rendering efficiency. Subsequent works have explored explicit grid-based representations (Müller et al., 2022; Fridovich-Keil et al., 2022; Chen et al., 2022; Sun et al., 2022) to enhance training efficiency. Nonetheless, these NeRF-based methods still face challenges of slow rendering due to dense sampling for each ray. In contrast, Kerbl et al. (2023) introduce 3D Gaussian Splatting, a novel explicit representation framework that employs a highly optimized custom CUDA rasterizer to achieve unparalleled rendering speeds with high-fidelity novel view synthesis for complex scenes.
@@
Neural Rendering for Dynamic Scenes. Synthesizing new views of dynamic scenes from a series of 2D images captured at different times presents a significant challenge. Recent advancements have extended NeRF to handle monocular or multi-object dynamic scenes by learning a mapping from spatio-temporal coordinates to color and density (Lombardi et al., 2019; Mildenhall et al., 2019; Pumarola et al., 2021; Li et al., 2022b; a; Cao & Johnson, 2023; Song et al., 2023; Attal et al., 2023; Fridovich-Keil et al., 2023; Wang et al., 2023). Unfortunately, these methods suffer from low rendering efficiency. To address this issue, some recent studies (Wu et al., 2024; Yang et al., 2024b; Das et al., 2024; Bae et al., 2024; Lu et al., 2024; Guo et al., 2024) have developed deformable 3D GS, which decouples dynamic scenes into a static canonical 3DGS and a deformation motion field to account for temporal variations in the 3D Gaussian parameters. Concurrently, a series of recent studies (Yang et al., 2024a; Duan et al., 2024; Li et al., 2024; Katsumata et al., 2024; Kratimenos et al., 2024) directly learn a set of spatio-temporal Gaussians to model static, dynamic, and transient content within a scene. However, these methods require a large number of Gaussians to achieve high-quality scene modeling, which brings expensive storage overhead. To this end, our work focuses on developing effective compression techniques for 4DGS (Yang et al., 2024a).
@@
3D Gaussian Splatting Compression. Since optimized scenes in 3DGS typically comprise millions of 3D Gaussians and require up to several gigabytes of storage, various compression strategies have been proposed to reduce the size, including redundant Gaussian pruning (Fan et al., 2024; Lee et al., 2024), spherical harmonics distillation or compactness (Lee et al., 2024; Fan et al., 2024; Niedermayr et al., 2024; Wang et al., 2024), vector quantization (Lee et al., 2024; Fan et al., 2024; Wang et al., 2024; Navaneet et al., 2024), and entropy models (Chen et al., 2024). However, due to the differences between 3DGS for static scene representation and 4DGS for dynamic scene representation, existing methods may be inapplicable to or unsuitable for 4DGS. In this paper, we aim to develop a more compact color representation and reduce the number of 4D Gaussians by considering temporal and viewpoint factors, thereby achieving a more efficient memory footprint. As far as we know, our study is among the first in 4DGS compression.
@@ 3 Method
In Section 3.1, we first review the technique of 4DGS (Yang et al., 2024a), which serves as the foundation of our method. Subsequently, in Section 3.2, we introduce how to develop our memory-efficient 4D Gaussian Splatting for modeling dynamic scenes. Finally, we detail the training process and describe how to store our compact 4DGS in Section 3.3.
@@ 3.1 Preliminary: 4D Gaussian Splatting
4D Gaussian Splatting (Yang et al., 2024a) optimizes a set of anisotropic 4D Gaussians via differentiable rasterization to effectively represent a dynamic scene. With a highly efficient rasterizer, the optimized model facilitates real-time rendering of high-fidelity novel views. Each 4D Gaussian is characterized by the following attributes: (i) 4D center $\bm{\mu}_{4D}=(\mu_{x},\mu_{y},\mu_{z},\mu_{t})^{T}\in\mathbb{R}^{4}$; (ii) 4D rotation $\bm{R}_{4D}$ represented by a pair of left quaternion $\bm{q}_{l}\in\mathbb{R}^{4}$ and right quaternion $\bm{q}_{r}\in\mathbb{R}^{4}$; (iii) 4D scaling factor $\bm{s}_{4D}=(s_{x},s_{y},s_{z},s_{t})^{T}\in\mathbb{R}^{4}$; (iv) time- and view-dependent RGB color represented by 4D spherical harmonics coefficients $\bm{h}\in\mathbb{R}^{3(k_{v}+1)^{2}(k_{t}+1)}$ with the view degrees of freedom $k_{v}$ and time degress of freedom $k_{t}$; (v) spatial opacity $o\in[0,1]$.
@@
Given 4D scaling matrix $S_{4D}=diag(\bm{s}_{4D})$ and 4D rotation matrix $\bm{R}_{4D}$, we parameterize 4D Gaussian’s covariance matrix as:
$\bm{\Sigma}_{4D}=\bm{R}_{4D}\bm{S}_{4D}\bm{S}_{4D}^{T}\bm{R}_{4D}^{T}=\begin{pmatrix}\mathbf{U}&\mathbf{V}\\ \mathbf{V}^{T}&\mathbf{W}\end{pmatrix},\mathbf{U}\in\mathbb{R}^{3\times 3}.$ (1)
@@
When rendering the scene at time $t$, each 4D Gaussian is sliced into 3D space. The density of the sliced 3D Gaussian at the spatial point $\bm{x}$ is expressed as:
$G_{3D}(\bm{x},t)=\sigma(t)e^{-\frac{1}{2}[\bm{x}-\bm{\mu}_{3D}(t)]^{T}\bm{\Sigma}^{-1}_{3D}[\bm{x}-\bm{\mu}_{3D}(t)]},$ (2)
@@
where $\bm{\Sigma}_{3D}=\mathbf{U}-\frac{\mathbf{V}\mathbf{V}^{T}}{\mathbf{W}}$ represents the time-invariant 3D covariance matrix. The temporal decay opacity, $\sigma(t)=e^{-\frac{(t-\mu_{t})^{2}}{2\mathbf{W}}}$, utilizes a 1D Gaussian function to modulate the contribution of each Gaussian to the $t$-th scene. The time-variant 3D center, $\bm{\mu}_{3D}(t)=\bm{\mu}_{3D}+(t-\mu_{t})\frac{\mathbf{V}}{\mathbf{W}}$, introduces a linear motion term to the 3D center position $\bm{\mu}_{3D}=(\mu_{x},\mu_{y},\mu_{z})^{T}$, assuming that all motions can be approximated as linear motion within a very small time range. After temporal slicing, the following process involves projecting sliced 3D Gaussians onto the 2D image plane based on depth order from specific view direction, and executing the fast differentiable rasterization to render the final image. Although this paradigm provides high-quality novel view synthesis, it necessitates large amount of Gaussians to fully reconstruct a dynamic scene, which brings unbearable storage overhead. This challenge drives our memory-efficient 4D Gaussian Splatting design.
Figure 3: Overview of our proposed memory-efficient Gaussian Splatting framework. (a) The original 4D Gaussian uses 4D spherical harmonics $\bm{h}$ to represent color, which is highly redundant and consumes substantial memory. (b) Our memory-efficient 4D Gaussian replaces $\bm{h}$ with a compact, view-independent, and time-independent color component $\bm{c}_{dc}$, achieving an about 8$\times$ reduction in storage overhead. (c) In the per-Gaussian transformation, a lightweight AC color predictor compensates for the absent viewpoint and temporal information in $\bm{c}_{dc}$, and a deformation predictor expands the action range of each Gaussian. (d) Our rendering process consists of four steps: per-Gaussian transformation, temporal slicing, projection, and differentiable rasterization.
3.2 Memory-efficient 4D Gaussian Splatting for Dynamic Scenes
Overview. As illustrated in Fig. 3, we develop our memory-efficient 4D Gaussian framework to significantly reduce the number of per-Gaussian stored parameters and drive the model to reconstruct dynamic scene with fewer 4D Gaussians. During the rendering process, we utilize a set of optimized 4D Gaussians and initially transform each Gaussian based on specific time and view direction. This transformation procedure involves Gaussian color prediction and geometry deformation. By modifying the geometric structure of each Gaussian, we effectively broaden its action range. This expansion not only reduces the total number of Gaussians required but also increases the rendering participation rate of each Gaussian. Following the per-Gaussian transformation, we adhere to the established protocols of the original 4DGS (Yang et al., 2024a) to carry out temporal slicing, projection, and differentiable rasterization, all critical for rendering high-quality frames.
Memory-efficient 4D Gaussian. 4DGS introduces 4D spherical harmonics $\bm{h}$ to model the temporal evolution of view-dependent color in dynamic scenes, which typically requires 144 of the total 161 parameters and contributes to the main storage overhead. While Lee et al. (2024) have explored the use of a grid-based neural field to replace SH coefficients $\bm{h}$, we find that directly applying this method results in severe performance loss compared to 4DGS (see Table 3).
To overcome this issue, we propose a compact DC-AC color (DAC) representation. Specifically, we decouple the color attribute as a per-Gaussian DC color component $\bm{c}_{dc}\in\mathbb{R}^{3}$ and a temporal-viewpoint aware AC color predictor $\mathcal{F}_{\bm{\phi}}$. To predict the final color $\bm{c}_{t,v}$ of each Gaussian, we first compute the normalized view direction $\bm{d}_{v}=\frac{\bm{\mu}_{3D}-\bm{p}_{v}}{||\bm{\mu}_{3D}-\bm{p}_{v}||_{2}}$ for each Gaussian according to the camera center point $\bm{p}_{v}\in\mathbb{R}^{3}$ at the viewpoint $v$. Then, we concatenate the 3D position $\bm{\mu}_{3D}$, view direction $\bm{d}_{i}$, time $t$, and DC color $\bm{c}_{dc}$ and input them to a lightweight MLP network $\mathcal{F}_{\bm{\phi}}$:
$\bm{c}_{t,v}={\rm sigmoid}(\bm{c}_{dc}+\mathcal{F}_{\bm{\phi}}({\rm sg}(\bm{\mu}_{3D}),{\rm sg}(\bm{d}_{v}),t,\bm{c}_{dc})),$ (3)
where ${\rm sg}(\cdot)$ indicates a stop-gradient operation. This hybrid color composition method not only effectively preserves the individual information using DC component and supplements the missing viewpoint and time information using the AC predictor to maintain high rendering quality, but also reduces the storage overhead by up to 8×8$\times$8 × compared to the original 4DGS (Yang et al., 2024a).

### v2

@@ Abstract
4D Gaussian Splatting (4DGS) has recently emerged as a promising technique for capturing complex dynamic 3D scenes with high fidelity. It utilizes a 4D Gaussian representation and a GPU-friendly rasterizer, enabling rapid rendering speeds. Despite its advantages, 4DGS faces significant challenges, notably the requirement of millions of 4D Gaussians, each with extensive associated attributes, leading to substantial memory and storage cost. This paper introduces a memory-efficient framework for 4DGS. We streamline the color attribute by decomposing it into a per-Gaussian direct color component with only 3 parameters and a shared lightweight alternating current color predictor. This approach eliminates the need for spherical harmonics coefficients, which typically involve up to 144 parameters in classic 4DGS, thereby creating a memory-efficient 4D Gaussian representation. Furthermore, we introduce an entropy-constrained Gaussian deformation technique that uses a deformation field to expand the action range of each Gaussian and integrates an opacity-based entropy loss to limit the number of Gaussians, thus forcing our model to use as few Gaussians as possible to fit a dynamic scene well. With simple half-precision storage and zip compression, our framework achieves a storage reduction by approximately 190$\times$ and 125$\times$ on the Technicolor and Neural 3D Video datasets, respectively, compared to the original 4DGS. Meanwhile, it maintains comparable rendering speeds and scene representation quality, setting a new standard in the field. Code is available at https://github.com/Xinjie-Q/MEGA.
Figure 1: Our approach significantly reduces storage requirements while maintaining comparable photorealistic quality and real-time rendering speed with 4D Gaussian Splatting (4DGS) [45]. The core idea is to develop a memory-efficient 4D Gaussian representation and use as few Gaussians as possible to fit dynamic scenes well. (a) 4DGS requires up to 13 million Gaussians to render the Birthday scene, whereas our method only needs 0.91 million Gaussians. (b) Quantitative comparisons of rendering quality, storage size, and speed against various competitive baselines on the Technicolor dataset.
@@ 1 Introduction
Dynamic scene reconstruction from multi-view videos is gaining widespread interest in computer vision and graphics due to its broad applications in virtual reality (VR), augmented reality (AR), and 3D content production. The emergence of neural radiance field (NeRF) [30] enables high-quality novel view synthesis from multi-view image inputs. It has been further extended to represent dynamic scenes by modeling a direct mapping from spatio-temporal coordinates to color and density [35, 22, 3]. Despite the impressive visual quality of NeRF-based methods, they require dense sampling along rays, leading to slow rendering speeds that hinder practical applications.
The recent introduction of 3D Gaussian Splatting (3DGS) [17] marks a significant shift in the field of novel view synthesis. This approach incorporates the explicit 3D Gaussian representation and differentiable tile-based rasterization to enable real-time rendering speeds that significantly outperform NeRF-based methods. Built on this framework, subsequent studies have developed 4D Gaussian Splatting (4DGS) [45, 10], which conceptualizes scene variations across different timestamps as 4D spatio-temporal Gaussian hyper-cylinder. As shown in Fig. 2, when depicting a 3D scene at a given timestamp, these 4D Gaussians will first be sliced into 3D Gaussians with time-varying positions and opacity. Then, the 3D Gaussians with the temporal decay opacity below a specific threshold are filtered out. This filtering operation helps 4DGS to describe the transient content such as emerging or vanishing objects. Finally, following 3DGS, the remaining 3D Gaussians are projected onto 2D screens through fast rasterization. By directly optimizing a collection of 4D Gaussians, 4DGS effectively captures both static and dynamic scene elements, thereby achieving photorealistic visual quality.
@@
Nevertheless, compacting the properties of the 4D Gaussian alone cannot effectively alleviate the problem of excessive number of Gaussians required. Existing 4DGS baselines [45, 10] assume that each sliced 4D Gaussian exhibits only linear movement over time while maintaining constant covariance, which means that the complex motion in the scene has to be modeled by a combination of multiple Gaussians. Moreover, as illustrated in Fig. 4 (a), only about 6% of Gaussians actively participate in rendering at any given time, because the temporal decay opacity forces each Gaussian to be visible only near its mean time center and invisible at other times. These inherent properties significantly limit the effective utilization of each Gaussian, thereby increasing the number of Gaussians needed for adequate scene rendering. To overcome this limitation, we introduce an efficient entropy-constrained Gaussian deformation field designed to expand the operational range of 4D Gaussians. This deformation model leverages both temporal and viewpoint information to accurately represent Gaussian motion, shape, and transience changes. Meanwhile, a spatial opacity-based entropy loss is introduced to push the spatial opacity of each Gaussian towards binary states (either one or zero). This adjustment aids in identifying and eliminating non-essential Gaussians that contribute minimally to the overall performance. In this way, our proposed strategy not only effectively reduces the number of Gaussians, but also improves the utilization rate of the Gaussians involved in rendering given the time and viewing angle. Finally, to store the parameters of our streamlined 4DGS, we employ 16-bit floating-point (FP16) precision with zip delta compression algorithm to achieve further reductions in memory footprint. In summary, our main contributions are three-fold:
@@ 2 Related Works
Neural Rendering for Static Scenes. Recently, the advent of neural rendering has attracted increasing interest in 3D scene representation and reconstruction. NeRF, pioneered by [30], represents the volume density and view-dependent emitted radiance of a 3D scene as a function of 5D coordinates (3D position and 2D viewing direction) using an MLP. However, the vanilla NeRF relies solely on a large MLP to store scene information, significantly limiting its training and rendering efficiency. Subsequent works have explored explicit grid-based representations [31, 12, 4, 39] to enhance training efficiency. Nonetheless, these NeRF-based methods still face challenges of slow rendering due to dense sampling for each ray. In contrast, [17] introduce 3D Gaussian Splatting, a novel explicit representation framework that employs a highly optimized custom CUDA rasterizer to achieve unparalleled rendering speeds with high-fidelity novel view synthesis for complex scenes. Subsequent studies have further improved the representation efficiency [26, 20, 15, 7, 47, 49] and have been extended to various vision understanding and editing applications [5, 37, 42, 48, 34].
@@
Neural Rendering for Dynamic Scenes. Synthesizing new views of dynamic scenes from a series of 2D images captured at different times presents a significant challenge. Recent advancements have extended NeRF to handle monocular or multi-object dynamic scenes by learning a mapping from spatio-temporal coordinates to color and density [25, 29, 35, 22, 21, 3, 38, 1, 13, 40, 24]. Unfortunately, these methods suffer from low rendering efficiency. To address this issue, some recent studies [43, 44, 9, 2, 27, 14] have developed deformable 3D GS, which decouples dynamic scenes into a static canonical 3DGS and a deformation motion field to account for temporal variations in the 3D Gaussian parameters. Concurrently, a series of recent studies [45, 10, 23, 16, 18] directly learn a set of spatio-temporal Gaussians to model static, dynamic, and transient content within a scene. However, these methods require a large number of Gaussians to achieve high-quality scene modeling, which brings expensive storage overhead. To this end, our work focuses on developing effective compression techniques for 4DGS [45].
@@
3D Gaussian Splatting Compression. Since optimized scenes in 3DGS typically comprise millions of 3D Gaussians and require up to several gigabytes of storage, various compression strategies have been proposed to reduce the size, including redundant Gaussian pruning [11, 20], spherical harmonics distillation or compactness [20, 11, 33, 41], vector quantization [20, 11, 41, 32], and entropy models [6]. However, due to the differences between 3DGS for static scene representation and 4DGS for dynamic scene representation, existing methods may be inapplicable to or unsuitable for 4DGS. In this paper, we aim to develop a more compact color representation and reduce the number of 4D Gaussians by considering temporal and viewpoint factors, thereby achieving a more efficient memory footprint.
Figure 3: Overview of our proposed memory-efficient Gaussian Splatting framework. (a) The original 4D Gaussian uses 4D spherical harmonics $\boldsymbol{h}$ to represent color, which is highly redundant and consumes substantial memory. (b) Our memory-efficient 4D Gaussian replaces $\boldsymbol{h}$ with a compact, view-independent, and time-independent color component $\boldsymbol{c}_{dc}$, achieving an about 8$\times$ reduction in storage overhead. (c) In the per-Gaussian transformation, a lightweight AC color predictor compensates for the absent viewpoint and temporal information in $\boldsymbol{c}_{dc}$, and a deformation predictor expands the action range of each Gaussian. (d) Our rendering process consists of four steps: per-Gaussian transformation, temporal slicing, projection, and differentiable rasterization.
@@ 3 Method
In Section 3.1, we first review the technique of 4DGS [45], which serves as the foundation of our method. Subsequently, in Section 3.2, we introduce how to develop our memory-efficient 4D Gaussian Splatting for modeling dynamic scenes. Finally, we detail the training process and describe how to store our compact 4DGS in Section 3.3.
@@ 3.1 Preliminary: 4D Gaussian Splatting
4D Gaussian Splatting [45] optimizes a set of anisotropic 4D Gaussians via differentiable rasterization to effectively represent a dynamic scene. With a highly efficient rasterizer, the optimized model facilitates real-time rendering of high-fidelity novel views. Each 4D Gaussian is characterized by the following attributes: (i) 4D center $\boldsymbol{\mu}_{4D}=(\mu_{x},\mu_{y},\mu_{z},\mu_{t})^{T}\in\mathbb{R}^{4}$; (ii) 4D rotation $\boldsymbol{R}_{4D}$ represented by a pair of left quaternion $\boldsymbol{q}_{l}\in\mathbb{R}^{4}$ and right quaternion $\boldsymbol{q}_{r}\in\mathbb{R}^{4}$; (iii) 4D scaling factor $\boldsymbol{s}_{4D}=(s_{x},s_{y},s_{z},s_{t})^{T}\in\mathbb{R}^{4}$; (iv) time- and view-dependent RGB color represented by 4D spherical harmonics coefficients $\boldsymbol{h}\in\mathbb{R}^{3(k_{v}+1)^{2}(k_{t}+1)}$ with the view degrees of freedom $k_{v}$ and time degress of freedom $k_{t}$; (v) spatial opacity $o\in[0,1]$.
@@
Given 4D scaling matrix $S_{4D}=diag(\boldsymbol{s}_{4D})$ and 4D rotation matrix $\boldsymbol{R}_{4D}$, we parameterize 4D Gaussian’s covariance matrix as:
$\boldsymbol{\Sigma}_{4D}=\boldsymbol{R}_{4D}\boldsymbol{S}_{4D}\boldsymbol{S}_{4D}^{T}\boldsymbol{R}_{4D}^{T}=\begin{pmatrix}\mathbf{U}&\mathbf{V}\\ \mathbf{V}^{T}&\mathbf{W}\end{pmatrix},$ (1)
@@
where $\mathbf{U}\in\mathbb{R}^{3\times 3}$. When rendering the scene at time $t$, each 4D Gaussian is sliced into 3D space. The density of the sliced 3D Gaussian at the spatial point $\boldsymbol{x}$ is expressed as:
$G_{3D}(\boldsymbol{x},t)=\sigma(t)e^{-\frac{1}{2}[\boldsymbol{x}-\boldsymbol{\mu}_{3D}(t)]^{T}\boldsymbol{\Sigma}^{-1}_{3D}[\boldsymbol{x}-\boldsymbol{\mu}_{3D}(t)]},$ (2)
@@
where $\boldsymbol{\Sigma}_{3D}=\mathbf{U}-\frac{\mathbf{V}\mathbf{V}^{T}}{\mathbf{W}}$ represents the time-invariant 3D covariance matrix. The temporal decay opacity, $\sigma(t)=e^{-\frac{(t-\mu_{t})^{2}}{2\mathbf{W}}}$, utilizes a 1D Gaussian function to modulate the contribution of each Gaussian to the $t$-th scene. The time-variant 3D center, $\boldsymbol{\mu}_{3D}(t)=\boldsymbol{\mu}_{3D}+(t-\mu_{t})\frac{\mathbf{V}}{\mathbf{W}}$, introduces a linear motion term to the 3D center position $\boldsymbol{\mu}_{3D}=(\mu_{x},\mu_{y},\mu_{z})^{T}$, assuming that all motions can be approximated as linear motion within a very small time range. After temporal slicing, the following process involves projecting sliced 3D Gaussians onto the 2D image plane based on depth order from specific view direction, and executing the fast differentiable rasterization to render the final image. Although this paradigm provides high-quality novel view synthesis, it necessitates large amount of Gaussians to fully reconstruct a dynamic scene, which brings unbearable storage overhead. This challenge drives our memory-efficient 4D Gaussian Splatting design.
3.2 Memory-efficient 4D Gaussian Splatting